OpenStudy (anonymous):

It is given that 10^x=3 and 10^y=7. What is the value of 10^x+y?

OpenStudy (anonymous):

Hint: 

OpenStudy (anonymous):

It's all in the properties of exponents! 3^2 X 3^3 = 3^(2+3) BECAUSE 9(3^2) X 27(3^3) = 234(which is 3^5)

OpenStudy (anonymous):

@mel2 can u do it?